OpenStudy (anonymous):

A right triangle has legs of length a and b, and hypotenuse of length c. Find the missing length. (a=the square root of 5, c=3.)

OpenStudy (saifoo.khan):

Use pythagoras!

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (saifoo.khan):

\[a^2 +b^2 + c^2 \]

OpenStudy (anonymous):

yes, but how does it work with a square root?

OpenStudy (anonymous):

(square root of 5)^2 = 5

OpenStudy (saifoo.khan):

\[(\sqrt{5})^2 + 3^2 = c^2 \] root cancels with square: \[\sqrt{5 + 9} = c\] c = \[\sqrt{14}\]

OpenStudy (anonymous):

oh ok thank you